# Toward Quantum Analogue Simulation of Many-Body Supersymmetry with Rydberg Atom Arrays

### v1

@@ Abstract
where $c_{N}$ is a numerical factor proportional to the slope of the degeneracy splitting $\delta$ with $\Delta$. To see the splitting we compare part of the spectrum at (away from) the SUSY point in Fig. 4a (Fig. 4b). Even though $\delta$ appears to be negligible, the change in $\widetilde{W}$ is around $10\%$ for $N=6$, as it depends on the slope determined by $c_{N}\beta$ ($c_{6}\approx 4\times 10^{-2}$). On the contrary, the variation in $\widetilde{W}$ for $N=9$ as a function of $\Delta$ is small. This can be attributed to a relatively small $c_{9}\approx 10^{-3}$.
Figure 4 (II) plots the case $N\neq 3j$ where the zero energy mode leads to topological protection of $\widetilde{W}$. We find $|\widetilde{W}-\widetilde{W}_{\rm SUSY}|$ to be negligibly small, suppressed by orders of magnitude compared to the $N=3j$ case, implying robustness. To illustrate this case, the energy spectrum at (Fig. 4d) and away (Fig. 4c) from the SUSY point is shown.
Figure 4c shows the zero energy mode shifted to a nonzero value. However, at low temperatures, this non-degenerate mode has nearly a unity occupation probability, and $\widetilde{W}$ remains insensitive to the breaking of the degeneracy of higher energy modes. The first order expansion is:
@@
where $E_{1}$ is the lowest $E>0$ state in the SUSY spectrum [51]. Comparing Eqs. (6) and  (7), we see the exponential factor ${\rm exp}(-\beta E_{1})$ in the latter. This term is responsible for the topological protection, thus explaining the negligible variation in $\widetilde{W}$ at low temperatures.
The lowest energy gap tends to decrease as the system sizes are increased. For fixed $\beta$, we therefore see that larger $N$ sectors, e.g., $N=10$ and $N=11$, have relatively larger deviations than the lower sectors [Fig. 4 (II) inset]. We have checked that our numerical results of $|\widetilde{W}-\widetilde{W}_{\rm SUSY}|$ are in good agreement with the analytical results in Eqs. (6) and (7) [51]. A similar topological protection of $\widetilde{W}$ manifests when $J$ is varied [51]. But we find that topological protection breaks down for temperatures near the gap, as is expected from Eq. (7) [51].
Summary and outlook- Symmetries encoded in supercharges can connect to observables, $F$, which, when processed with appropriate averaging protocols, reveals SUSY. We proposed a normalized Witten index as an observable and have studied it in an XXZ spin model relevant for ongoing experiments with Rydberg atom arrays. Through our open-system numerical simulations we have demonstrated observability of the normalized Witten index and topological protection arising from zero energy modes.

### v2

@@ Abstract
where $c_{N}$ is a numerical factor proportional to the slope of the degeneracy splitting $\delta_{N}$ with $\Delta$. To see the splitting we compare part of the spectrum at (away from) the SUSY point in Fig. 4(a) [Fig. 4(b)]. See Ref. 48 for discussion over finite size effects and degeneracy splitting.
Figure 4 (II) plots the case $N\neq 3j$ where the zero energy mode leads to topological protection of $\widetilde{W}$. We find $|\widetilde{W}-\widetilde{W}_{\rm SUSY}|$ to be suppressed by orders of magnitude compared to the $N=3j$ case, implying robustness. To illustrate this case, the energy spectrum at [Fig. 4(c)] and away [Fig. 4(d)] from the SUSY point is shown.
Figure 4(d) shows the zero energy mode shifted to a nonzero value. However, at low temperatures, this nondegenerate mode has nearly a unity probability, and $\widetilde{W}$ remains insensitive to the breaking of the degeneracy of higher energy modes. The first order expansion is--
@@
where $E_{1}$ is the lowest $E>0$ state in the SUSY spectrum [48]. Comparing Eqs. (6) and  (7), we see that the factor ${\rm exp}(-\beta E_{1})$ is responsible for the vanishing slope and thus the topological protection, thereby explaining the negligible variation in $\widetilde{W}$ at low temperatures.
Our numerical results of $|\widetilde{W}-\widetilde{W}_{\rm SUSY}|$ are in good agreement with the analytical results in Eqs. (6) and (7) [48]. A similar topological protection of $\widetilde{W}$ manifests when $J$ is varied [48]. But we find that topological protection breaks down for temperatures near the gap, as is expected from Eq. (7) [48]. This topological protection demonstrates that the SUSY is robust against fluctuations provided they are smaller than the gap.
Summary and outlook-- Symmetries encoded in supercharges can connect to observables, $\hat{F}$, which, when processed with appropriate averaging protocols, reveals SUSY. We proposed a normalized Witten index as an observable and have constructed a corresponding averaging protocol. We studied the protocol in an XXZ spin model relevant for ongoing experiments with Rydberg atom arrays. Through our open-system numerical simulations we have demonstrated observability of the normalized Witten index and topological protection arising from zero energy modes.